Suppose you need to have $50,806.00 in an account 22.00 years from today and that the account pays 15.00%. How much do you have to deposit into the account 4.00 years from today?

Please help me how to input into BA II Plus Calculator and solving it out---Thanks
Answer #1

Number of years of Compounding = 22-4 = 18

Rate of interest = 15%

Amount required to be deposited i.e. present value = Future amount required/(1+Interest rate)^number of years

= 50,806/(1.15)^18

= $4,105.3849

i.e. $4,105.38